OpenStudy (anonymous):

How many 6 letter words (real or imaginary) can be made from the word MESSAGE? Please write the formula down for me, and the steps of how you got the answer.

OpenStudy (anonymous):

7x6x5x4x3x2x1 = 5040 possible outcomes.

OpenStudy (anonymous):

Because there are 7 letters in MASSAGE so this is the equation you would use.

OpenStudy (anonymous):

havent you gotta take into account that u have two e's and 2 ss's - im not sure

OpenStudy (anonymous):

thats wrong the answer in my text book says 360 words but i wanta kno how they did it. cuz ima be tested on it 2marro. and idk how 2 do it

OpenStudy (anonymous):

i have a pic. of another similar problem in the text book, but i need help understanding there example. do u want me to attach the pic so u could c the example

OpenStudy (anonymous):

6x5x4x3 is the right answer, giving you 360. I looked at MASSAGE because it's a 7 letter word, but I forgot we were solving for 6 letter words.

OpenStudy (anonymous):

so how u got that?

OpenStudy (anonymous):

Confirmed that is the right answer

OpenStudy (anonymous):

vermont is it massage or message?

OpenStudy (anonymous):

It is MASSAGE.

OpenStudy (sasogeek):

in combinations, the formula used to compute ur answer is n!/(n-r)! n=7 r=6 7 combination 6 = 7!/(7-6)! = 7! = 7x6x5x4x3x2x1 = 5040

OpenStudy (anonymous):

That's originally what I came up with sasogeek, but read up ^.

OpenStudy (sasogeek):

oh, message has 2 e's lol, forgot about that :P thanks for the heads up ;)

OpenStudy (anonymous):

im pretty rusty at these bu i think u must take in to account the fact that there are 2 A's and 2 S's

OpenStudy (sasogeek):

^

OpenStudy (anonymous):

Oh wow it is MESSAGE forgive me.

OpenStudy (anonymous):

lol - i thought it was

OpenStudy (anonymous):

So my final answer stands, 360 (6x5x4x3x2x1).

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so could u explain how u got ur answer?

OpenStudy (sasogeek):

I think it is MASSAGE because if you had MESSAGE, you have 2 E's and 2 S's, then you would consider them as 1 and that would mean 5 letters... Vermont

OpenStudy (sasogeek):

you cannot compute 5 combination 6 :/

OpenStudy (anonymous):

hey can some1 that understands the problem explain the formula they used 2 get the answer. plez and thanks

OpenStudy (anonymous):

the way i see it is as its 6 from 7 number of possibiliites is 6x5x4x3x2x1 but as there are a doble number you divide by 2 to give 6x5x4x3 = 360

OpenStudy (anonymous):

not sure about that though

OpenStudy (anonymous):

yea but when u did 6x5x4x3 what happened to the 2 and 1 at the end, why does it stop at 3??

OpenStudy (anonymous):

^

OpenStudy (anonymous):

i explained about the 2 - because 6 letter word has to containe two duplicate letters

OpenStudy (sasogeek):

but if you had MASSAGE, you have only 2 S's and hence 6!/(6-6)!*2! = 6*5*4*3*2*1/(1)*2*1 the 2*1 below cancels the one on top so u end at 3 =6*5*4*3 = 360

OpenStudy (sasogeek):

dang! message has 2 e's, massage has 2 a's, i guess it all works out the same way

OpenStudy (sasogeek):

not sure if that is still right but we still end up with 360 so :|

OpenStudy (anonymous):

i'm about 99% sure thats correct as i've not done this sort of stuff for yearrssss